OpenStudy (anonymous):

Calculus Related Rates: If a snowball melts so that its surface area decreases at a rate of 7 cm2/min, find the rate at which the diameter decreases when the diameter is 10 cm. Please show all of your work... and tell me why I keep getting -.0139?

OpenStudy (anonymous):

SA=4pi r^2 or 4pi (1/2D)^2

OpenStudy (anonymous):

\[S(D)= \pi D^2\] so \[S'=2\pi D D'\]

OpenStudy (anonymous):

you are told that \[S'=-7\] so solve for \[D'\] and you have it

OpenStudy (ybarrap):

A =Surface Area of a Sphere = 4pi*r^2 dA/dt = - 7cm^2/min dA/dt = dA/dr * dr/dt, where dr/dt is change in radius wrt time dA/dt = 8 * pi *r * dr/dt = -7 cm^2/min so dr/dt = -7/(pi*r*8) = -7/pi*5*8) = -.056 cm/min Ans. -.056 cm/min

OpenStudy (anonymous):

dA / dt = dA/ dD * dD/dt -7 = 2*pi* D * dD / dt

myininaya (myininaya):

\[SA=4 \pi r^2 => SA'=4 \pi (2 r r')=8 \pi r r' \] where SA is the surface area \[SA'=-7\] \[d=10 => r=\frac{10}{2}=5\] we need to find d' d=2r => d'=2r' so we have \[-7=8 \pi (5) r' => r'=\frac{-7}{8 \pi (5)}=\frac{-7}{40 \pi} => d'=2r'=2(\frac{-7}{40 \pi})=\frac{-7}{20 \pi}\]

OpenStudy (anonymous):

ybarrap- Your answer would be in terms of r right, not D? And give me a second to work this through now, and I will let you all know if Webassign agrees.

OpenStudy (anonymous):

Okay thanks everyone. The answer was actually positive because it asked for the rate in which it decreases. It was 7/20pi

OpenStudy (anonymous):

i get rate of change = -7/20 pi - same as myin - thatgs rate of change of diameter at d = 10 cm